A car is traveling towards the east at a speed of 10.0 ms1. It then turns towards the

north and continues to travel at the same speed. If the turn is made in a time of 5.0 s and

the mass of the car is 2000 kg, calculate (i) the impulse delivered to the car because of

the turn and (ii) the average force exerted on the car during the turn.


1
Expert's answer
2021-03-23T16:57:37-0400
"P=m\times v"

"v=v_0+at"

"a=\frac{v^2}{R}"

"\frac{\pi R}{2}=vt"

"R=32 m"

"v=25.6"

"P=2000\times25.6=51,200 N"

"F=ma"

"F=2000\times20.48=40,960N"